How much acid must be added to 60 grams of a solution that is 65% acid to obtain a new solution that is 75% acid?
Answer by checkley77(12844)

.65*60+x=.75(60+x)
39+x=45+.75x
x-.75x=45-39
.25x=6
x=6/.25
x=24 grams of acid is used.
Proof:
.65*60+24=.75(60+24)
39+24=.75*84
63=63